How to find the answer for root over 8+3i ?????????????????

Grade:11

1 Answers

Riddhish Bhalodia
8 years ago
we assume the root of the complex number be a + ib
then we have
\sqrt{8+3i} = a+ib
square and solve for a and b by comaring real and imaginary parts